A total of twelve thousand passengers normally ride the green line of the MBTA during te morning rush hour. The token prices for a ride are .25 for childern under 12, $1 for adults and .50 for senior citizens. The revenue from these riders is $10,700. If he token prices were raised to .35 for childern under 12 and $1.50 for adults, and the senior ctizens price was unchanged, the expected revenue from these riders would be $15,820. How many riders in each category normally ride the green line during the morning rush hour?
I know how to do matrices, however, I can not figure out how to set this one up and could use some help.

Answer by stanbon(75887)

A total of twelve thousand passengers normally ride the green line of the MBTA during te morning rush hour.
a + s + c = 12000
-------------------------
The token prices for a ride are .25 for childern under 12, $1 for adults and .50 for senior citizens. The revenue from these riders is $10,700.
100a + 50s + 35c = 1070000 cents
-------------------------
If the token prices were raised to .35 for childern under 12 and $1.50 for adults, and the senior ctizens price was unchanged, the expected revenue from these riders would be $15,820.
150a + 50s + 35c = 1582000 cents
-------------------------
System to solve:
a + s + c = 12000
100a + 50s + 35c = 1070000
150a + 50s + 35c = 1582000
-------------------------
Use what you know about matrices to solve for a,s, and c.
==========================================
How many riders in each category normally ride the green line during the morning rush hour?
---------------------------------------------------------------------
Cheers,
Stan H.